the officers of science club are planning to sell 125 tickets to be raffled during the school's foundation day. One ticket will win Php2,000 and the other tickets will win nothing. If you will buy one ticket, what will be your expected gain E(X)?


1
Expert's answer
2021-05-24T19:14:50-0400

Probability of winning Php2000 = "\dfrac{1}{125}=0.008"


Probability of not winning any money = "1-\dfrac{1}{125}=1-0.008=0.992"


So, table will be:



and we know that,

Expected gain "E(X)=\sum X\cdot P(X)"

"E(X)= 0+16=16"


So, If you will buy one ticket, your expected gain "\boxed{E(X)=16}"